# Virasoro Symmetry in Neural Network Field Theories

### v1

@@ Abstract
Introduction.— The interface between theoretical physics and machine learning Carleo:2019ptp has revealed a deep correspondence between Neural Networks (NNs) and Quantum Field Theory (QFT). In the infinite width limit, the output of a randomly initialized network converges to a Gaussian Process governed by a kernel $K(x,y)$ Neal:1996 ; NIPS1996_ae5e3ce4 ; Lee:2017wiy , effectively defining a free field theory Halverson:2020trp ; Halverson:2021aot . Finite-width corrections induce non-Gaussian interactions, providing a framework to simulate renormalization group flows and neural effective actions Dyer:2019uzd ; Yaida:2020 ; Roberts:2021principles ; Erdmenger:2022tqi ; Grosvenor:2022edge ; Schoenholz:2016 .
Central to the development of this program is building NNs that encode spacetime symmetries. While global conformal invariance ($SO(d+1,1)$) has been achieved via embedding space formalism Halverson:2024axc and extended to include conformal defects Billo:2016cpy ; Capuozzo:2025ozt , these constructions typically yield GFFs. While respecting global conformal symmetries, for generic scaling dimensions, these GFFs lack a local, conserved stress-energy tensor $T_{\mu\nu}$ satisfying conformal Ward identities. For NN-FTs with arbitrary dimensional input space, this would present an obstruction to the standard computation of Weyl anomalies but would still allow for the computation of correlation functions and OPEs. In two dimensions (2d), though, this is a fatal deficit: without a local $T(z)$, we cannot construct Virasoro generators $L_{n}$, and thus no local conformal symmetry. The lack of local conformal symmetry in 2d presents a significant barrier to the application of the NN-FT formalism to the study of, e.g., low dimensional critical phenomena and string theory.
@@
where phases $\gamma_{j}$ are drawn uniformly from $0,2\pi)$ and amplitudes $A_{j}=1$ are fixed. Here we note that any single realization of the network $\phi(x;\Theta)$ does not respect translation invariance, the ensemble does. The uniform integration over phases $\gamma_{j}$ enforces functional dependence only on separations, e.g. $\langle\phi(x)\phi(y)\rangle=f(x-y)$. Furthermore, although the Log-correlated field $\phi(x)$ shifts under special conformal transformations, its gradients $\partial\phi$ and Neural Vertex Operators (NVOs) $\mathcal{V}_{\alpha}$ transform covariantly. The crucial ingredient to restoring global conformal invariance is the distribution of wavevectors $\vec{k}_{j}$. To ensure scale invariance of the Gaussian Process kernel, the probability measure $d\mu(k)=p(k)d^{2}k$ must be invariant under dilatations $k\to\lambda k$. This fixes the spectral density to be $p(k)\propto|k|^{-2}$ in the infinite width limit.
@@
Discussion.— We have established a framework for constructing local 2d CFTs from NNs, identifying the LK architecture as a scalar Gaussian fixed point with Virasoro symmetry. By utilizing non-linear readouts and the method of images, we further demonstrated that NN-FTs can simulate theories with fermionic statistics and boundaries.

### v2

@@ Abstract
Introduction.— The interface between theoretical physics and machine learning Carleo:2019ptp has revealed a deep correspondence between Neural Networks (NNs) and Quantum Field Theory (QFT). In the infinite width limit, the output of a randomly initialized network converges to a Gaussian Process governed by a kernel $K(x,y)$ Neal:1996 ; NIPS1996_ae5e3ce4 ; Lee:2017wiy , effectively defining a free field theory Halverson:2020trp ; Halverson:2021aot . Finite-width corrections induce non-Gaussian interaction terms, providing a framework to simulate renormalization group flows and neural effective actions Dyer:2019uzd ; Yaida:2020 ; Demirtas:2023fir ; Roberts:2021principles ; Erdmenger:2022tqi ; Grosvenor:2022edge ; Schoenholz:2016 .
Central to the development of this program is building NNs that encode symmetries Maiti:2021fpy . While global conformal symmetry ($SO(d+1,1)$) has been achieved via embedding space formalism Halverson:2024axc and extended to include conformal defects Billo:2016cpy ; Capuozzo:2025ozt , these constructions typically yield GFFs. While respecting global conformal symmetries, for generic scaling dimensions, these GFFs lack a local, conserved stress-energy tensor $T_{\mu\nu}$ satisfying conformal Ward identities. For NN-FTs with arbitrary dimensional input space, this would present an obstruction to the standard computation of Weyl anomalies but would still allow for the computation of correlation functions and OPEs. In two dimensions (2d), though, this is a fatal deficit: without a local $T(z)$, we cannot construct Virasoro generators $L_{n}$, and thus no local conformal symmetry. The lack of local conformal symmetry in 2d presents a significant barrier to the application of the NN-FT formalism to the study of, e.g., low dimensional critical phenomena and string theory.
@@
where phases $\gamma_{j}$ are drawn uniformly from $0,2\pi)$ and amplitudes $A_{j}=1$ are fixed. Unlike in Halverson:2024axc where the $d$-dimensional neural free boson and its stress tensor were first derived, we note that in the cos-net architecture any single realization of the network $\phi(x)$ does not manifestly respect translation invariance; rather, the ensemble does. The uniform integration over phases $\gamma_{j}$ enforces functional dependence only on separations, e.g. $\langle\phi(x)\phi(y)\rangle=f(x-y)$. Furthermore, although $\phi(x)$ shifts under special conformal transformations, its gradients $\partial\phi$ and Neural Vertex Operators (NVOs) $\mathcal{V}_{\alpha}$ transform covariantly. Crucially, conformal invariance of the correlators is realized through the distribution of wavevectors $\vec{k}_{j}$. To ensure scale invariance of the Gaussian Process kernel, the probability measure $d\mu(k)=p(k)d^{2}k$ must be invariant under dilatations $k\to\lambda k$. This fixes the spectral density to be $p(k)\propto|k|^{-2}$ in the infinite width limit.
@@
Discussion.— We have established a framework for constructing local 2d CFTs from NNs, identifying the “Log-Kernel” architecture as a scalar Gaussian fixed point with Virasoro symmetry. By utilizing non-linear readouts and the method of images, we further demonstrated that NN-FTs can simulate compact target spaces, fermionic statistics, and boundaries.
@@ B.2.2 Ramond (R) Sector and the Zero Mode
B.3 Ghosts in the Machine
Here we note that the above constructions can easily be extended to neural ghost fields. That is, we have the architectures to build fields with ‘bosonic’ or ‘fermionic’ statistics. We demonstrate below that by choosing the spectral prior along with Gaussian or Grassmann valued output weights, we can define neural bc and $\beta\gamma$ ghosts.
B.3.1 $bc$ ghosts
The fermionic $bc$-system consists of two anti-commuting network outputs $b(z)$ and $c(z)$ with scaling dimensions $h_{b}=\lambda$ and $h_{c}=1-\lambda$. We realize this system using an ensemble of $N$ neurons with spatial frequencies $\vec{k}_{j}\sim p(k)\propto 1/|k|$ and Grassmann-valued output weights $\xi_{j,1},\xi_{j,2}$. The network outputs are defined as:
$\begin{split}b(z;\Theta)&=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left(\xi_{j,1}u_{j}(z)+i\xi_{j,2}u_{j}(z)\right)~,\\ c(z;\Theta)&=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left(\xi_{j,1}v_{j}(z)-i\xi_{j,2}v_{j}(z)\right)~,\end{split}$ (S11)
where $u_{j},v_{j}$ are the holomorphic basis functions
$\begin{split}u_{j}(z)&=\sqrt{2k_{j}}e^{-i\theta_{k_{j}}}\cos(k_{j}\cdot z)~,\\ v_{j}(z)&=i\sqrt{2k_{j}}e^{-i\theta_{k_{j}}}\sin(k_{j}\cdot z)~.\end{split}$ (S12)
This architecture enforces the propagator $\mathbb{E}[b(z)c(w)]=(z-w)^{-1}$.
The geometry of the emergent field theory is defined by how we measure the “stress-energy tensor” of the network state. We construct a specific readout head $T^{(\lambda)}$ that combines the network outputs and their gradients. We introduce a tunable readout twist hyperparameter $\lambda$ such that the readout head is defined as
$T^{(\lambda)}(z;\Theta)=:(\partial_{z}b(z;\Theta))c(z;\Theta):-\lambda\partial_{z}(:b(z;\Theta)c(z;\Theta):)~.$ (S13)
Here, $\partial_{z}$ acts analytically on the basis functions $u_{j}(z),v_{j}(z)$. The notation :AB::AB: implies subtraction of the initialization variance (vacuum energy):
$A(z)B(z):\equiv A(z)B(z)-\mathbb{E}_{\Theta}[A(z)B(z)]~.$ (S14)
The central charge $c$ is a measure of the variance of the stress-energy tensor readout across the random initialization ensemble.
$\text{Cov}(T(z),T(w))=\mathbb{E}_{\Theta}[T^{(\lambda)}(z)T^{(\lambda)}(w)]_{\text{connected}}~.$ (S15)
Let $A=(1-\lambda)$ and $B=-\lambda$. The stress tensor variable is:
$T(\Theta)=A(\partial b)c+Bb(\partial c)~.$ (S16)
We evaluate the ensemble expectation using Isserlis’ Theorem for Fermions (Wick’s Theorem). The variance decomposes into integrals over the kernel derivatives $\partial K(z,w)$:
$\mathbb{E}[T(z)T(w)]=\mathbb{E}[(\partial bc)_{z}(\partial bc)_{w}]+\mathbb{E}[(b\partial c)_{z}(b\partial c)_{w}]-2\mathbb{E}[(\partial b\partial c)_{z}(bc)_{w}]~.$ (S17)
Substituting the kernel $K(z,w)=(z-w)^{-1}$, the first term gives
$A^{2}\mathbb{E}[\partial b(z)c(w)]\mathbb{E}[c(z)\partial b(w)]=-A^{2}(\partial_{z}K)(\partial_{w}K)=\frac{A^{2}}{z_{12}^{4}}~.$ (S18)
The second term evaluates to
$B^{2}\mathbb{E}[b(z)\partial c(w)]\mathbb{E}[\partial c(z)b(w)]=-B^{2}(\partial_{w}K)(\partial_{z}K)=\frac{B^{2}}{z_{12}^{4}}~,$ (S19)
and the final term becomes
$2\mathbb{E}[\partial b(z)\partial c(w)]\mathbb{E}[c(z)b(w)]=4AB\frac{1}{z_{12}^{4}}~.$ (S20)
The total variance of (twisted) stress-energy tensor readouts determines the central charge for the neural $bc$-ghosts to
$\begin{split}c&=2(4\lambda(1-\lambda)-(1-\lambda)^{2}-\lambda^{2})=1-3(2\lambda-1)^{2}~,\end{split}$ (S21)
where the overall sign comes from the fermionic statistics. Substituting $\lambda=2$, we get the expected critical value
$c_{bc}=1-3(2(2)-1)^{2}=-26~.$ (S22)
B.3.2 $\beta\gamma$ ghosts
The $\beta\gamma$ ghosts are engineered with commuting statistics; hence realized as symplectic bosons. We utilize the same $1/k$ spectral prior but sample output weights from independent Gaussian distributions $w_{j,a}\sim\mathcal{N}(0,1)$. To reproduce the first-order propagator, we organize the readout into symplectic pairs:
$\gamma(z;\Theta)$ $=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left(w_{j,1}u_{j}(z)+w_{j,2}v_{j}(z)\right)~,$ (S23)
$\beta(z;\Theta)$ $=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left(w_{j,1}v_{j}(z)-w_{j,2}u_{j}(z)\right)~.$ (S24)
The sign inversion in $\beta$ encodes the symplectic structure.
The cross-correlation computes the symplectic product of the basis functions:
$\mathbb{E}_{\Theta}[\beta(z)\gamma(w)]$ $=\frac{1}{N}\sum_{j}\mathbb{E}\left[(w_{j,1}v_{j}(z)-w_{j,2}u_{j}(z))(w_{j,1}u_{j}(w)+w_{j,2}v_{j}(w))\right]$
$=\frac{1}{N}\sum_{j}(v_{j}(z)u_{j}(w)-u_{j}(z)v_{j}(w))$
$\underset{N\to\infty}{\longrightarrow}\int dk\,p(k)\left(v(z)u(w)-u(z)v(w)\right)$ (S25)
Going from the first to second line, we used the weight statistics $\mathbb{E}_{\Theta}[w_{j,a}w_{l,b}]=\delta_{jl}\delta_{ab}$ to eliminate the cross term and collapse the double sum in the first two terms. Substituting the definition of the mode functions and integrating against the spectral prior reproduces the Cauchy kernel:
$\mathbb{E}_{\Theta}[\beta(z)\gamma(w)]=-\frac{1}{z-w}~.$ (S26)
The stress tensor readout is defined in the same was as the $bc$ ghosts system above
$T^{(\lambda)}(z)=:(\partial\beta)\gamma:-\lambda\partial(:\beta\gamma:)~.$ (S27)
However, the weights $w$ are bosonic. When computing the variance $\mathbb{E}[TT]$, the closed loops do not acquire a minus sign. The resulting variance determines the central charge
$c=2(A^{2}+B^{2}-4AB)=3(2\lambda-1)^{2}-1$ (S28)
Substituting $\lambda=3/2$ gives
$c_{{\rm crit},\beta\gamma}=11$ (S29)
which recovers the expected critical value.